### v1

# A Simple Framework to Accelerate Multilingual Language Model for Monolingual Text Generation

Abstract
Recent advancements in large language models have facilitated the execution of complex language tasks, not only in English but also in non-English languages. However, the tokenizers of most language models, such as Llama, trained on English-centric corpora, tend to excessively fragment tokens in non-English languages. This issue is especially pronounced in non-roman alphabetic languages, which are often divided at a character or even Unicode level, leading to slower text generation. To address this, our study introduces a novel framework designed to expedite text generation in these languages. This framework predicts larger linguistic units than those of conventional multilingual tokenizers and is specifically tailored to the target language, thereby reducing the number of decoding steps required. Our empirical results demonstrate that the proposed framework increases the generation speed by a factor of 1.9 compared to standard decoding while maintaining the performance of a pre-trained multilingual model on monolingual tasks.
1 Introduction
The advent of large language models (LLMs) (OpenAI, 2023; Touvron et al., 2023; Antropic, 2023) has revolutionized the field of generative artificial intelligence, demonstrating remarkable capabilities to respond across diverse tasks and languages (Eloundou et al., 2023; Solaiman et al., 2023). These models, despite being predominantly trained on English data, exhibit significant multilingual proficiency (Bandarkar et al., 2023).
However, when applied to non-alphabetic languages, these models often suffer from slower text generation due to English-centric tokenization  Rust et al. (2021); Ahia et al. (2023); Petrov et al. (2023). Current tokenization techniques used in Large Language Models (LLMs) are data-driven and optimize segmentation based on the frequency of characters or bytes within a specific corpus (Sennrich et al., 2016; Kudo, 2018). As a result, the tokenizers of multilingual models, which are heavily influenced by English-dominant training data, are predominantly composed of English subwords. This leads to excessive fragmentation, where non-English words are overly segmented into a large number of subword units Rust et al. (2021); Ahia et al. (2023); Petrov et al. (2023). The autoregressive nature of LLMs further amplifies this inefficiency, as it sequentially requires the generation of text.
Figure 1: Overview of the proposed framework. Illustration of (Left) the generation with a pre-trained multilingual model and (Right) the generation of MuMo Framework. Given the Korean prefix “천왕성은” (Uranus is), the model generates the consecutive phrase “태양으로부터”(from the Sun) that consisted of 3 morphemes (“태양”, “으로”, “부터”) in Korean. The generation with the pre-trained multilingual model faces inefficiency due to excessive fragmentation, requiring 12 steps to generate only 3 Korean morphemes. However, the MuMo framework empowers the multilingual language model to generate multiple tokens in a single iteration by extracting a word from the Korean Vocabulary, requiring 3 steps.
To address these challenges, previous studies Wang et al. (2019b); Rust et al. (2021); Cui et al. (2023) have proposed replacing or augmenting the existing vocabulary of pre-trained multilingual models with language-specific vocabularies to more effectively encode monolingual text corpora. Specifically, Rust et al. (2021) improved mBERT Devlin et al. (2019) by replacing its tokenizer with a monolingual one and incorporating an additional 100,000 pre-training steps. On the other hand, Cui et al. (2023) enhanced Llama (Touvron et al., 2023) by expanding the Chinese vocabulary and further pre-training it on a 120GB text corpus that includes Chinese texts. However, this approach requires an extensive pre-training phase with a substantial amount of data.
Another approach to address the challenges is the use of small draft models (Leviathan et al., 2023; Chen et al., 2023a). These models generate draft output tokens, which are then verified by the original language model. However, a significant challenge arises when trying to identify or train a suitable small model that can handle multiple languages with reliable performance (Conneau et al., 2020; Bandarkar et al., 2023).
In response to these challenges, our research introduces MuMo, a novel framework designed to accelerate Multilingual language models for targeted Monolingual text generation, particularly in non-alphabetic languages. MuMo incorporates a new vocabulary of a target language into the output layer, also known as the Language Model (LM) head, and predicts the next token from this expanded vocabulary. This approach requires training only the extended portion of the output layer and specific layers of the feed-forward network. Importantly, MuMo eliminates the need for extensive text corpora or a draft model, requiring only a modest corpus of the target language, approximately 365M tokens in size.
Empirical results across summarization, and translation tasks in Korean and Japanese demonstrate that MuMo significantly accelerates text generation, achieving over a 1.9-fold increase in speed without compromising output quality.
The contributions of our research are as follows:
  • We propose MuMo, a framework that accelerates the decoding speed of non-Roman alphabetic languages without compromising the performance of the language model.
  • Our study shows a novel approach to address the issue of slow inference speed of non-English languages, a problem that arises due to the excessive fragmentation in pre-trained models that are primarily English-centric.
  • It can be effectively employed even in the absence of extensive computing resources or large monolingual text corpora.
2 Background
Language Modeling
The goal of auto-regressive language modeling is to create a probability distribution, denoted as $p_{\theta}(\mathbf{x})$, where $\theta$ represents the model parameter. This distribution is established over a sequence of tokens, represented as $\mathbf{x}=[x_{1},\ldots,x_{T}]$. These tokens are derived from a pre-established vocabulary, denoted as $\mathcal{V}$, where each token $x_{t}$ is a member of $\mathcal{V}$. A conventional approach involves maximizing the log-likelihood of a sequence:
$\max_{\theta}\mathcal{L}_{\text{MLE}}(p_{\theta},\mathbf{x})=\sum_{t=1}^{T}\log p_{\theta}(x_{t}|\mathbf{x}_{<t}).$ (1)
Tokenization in LM
Tokenization is the process of segmenting text into subword chunks (Kudo, 2018; Sennrich et al., 2016). The process can be started with a base vocabulary composed solely of characters, then, new vocabulary items are subsequently added by recursively merging existing ones based on the frequency statistics within a training corpus. However, in a multilingual setup, even a base vocabulary composed solely of characters can be extensive, with over 130K items (Wang et al., 2019a).
Byte-level Subword Tokenization
Radford et al. (2019) proposed using bytes as the basic units, with only 256 tokens required to form subwords. This Byte-level Byte Pair Encoding (BBPE) does not suffer from out-of-vocabulary problem and has become the standard in most pre-trained models (Brown et al., 2020; Muennighoff et al., 2022; Zhang et al., 2022b; Scao et al., 2022). However, BBPE often produces longer sequence length when handling multilingual scripts compared to character-based BPE (Wang et al., 2019a; Petrov et al., 2023).
Lang Word Multilingual Tokens Ko 서울 (“서”, “\0xec”, “\0xb8”, “\0x9a”) Ja 発売 (“発”, “\0xe5”, “\0xa3”, “\0xb2
Table 1: Examples of the tokenization results. These examples are tokenized by the Llama tokenizer. The target monolingual word are excessively segmented into byte units, when a suitable match is not found in the multilingual vocabulary.
Exccessive Fragmentation
We explore the issue of excessive fragmentation (Ahia et al., 2023; Petrov et al., 2023) in multilingual setups, a common challenge in multilingual tokenizers (Rust et al., 2021; Zhang et al., 2022a; Ahia et al., 2023; Petrov et al., 2023). The problem occurs when a tokenizer produces too many tokens due to significant differences between input texts and the pre-trained corpus (Rust et al., 2021; Hong et al., 2021).
The difficulty of integrating numerous languages into one model while balancing vocabulary composition contributes to this issue (Wang et al., 2019b). As the number of vocabulary items increases, model parameters expand exponentially, impacting training efficiency (Liang et al., 2023).
To investigate this, we compared the tokenization outputs of several multilingual model tokenizers: Bloomz (Scao et al., 2022), Llama-2 (Touvron et al., 2023), and GPT-3.5/4 (OpenAI, 2023). We used the FLoRes-200 (Goyal et al., 2022) dataset, which includes Wikipedia articles in 200 languages. Table 2 indicates that multilingual tokenizers generate more tokens for Japanese and Korean languages than their monolingual counterparts.
3 Proposed Framework
Figure 2: Illustration of a single step prediction with MuMo. Initially, the MuMo LM Head $f_{\text{mumo}}$ selects the top 6 candidates. Then, the pre-trained multilingual model verifies the feasibility of the candidates. Among the modules in MuMo, the Target Monolingual LM head (the Korean LM Head in the figure) is only trained.
We propose a framework named MuMo to accelerate the inference speed of a pre-trained multilingual LM for a non-alphabetic monolingual language via a given small monolingual dataset. In the section, we introduce 1) the model architecture, 2) the fine-tuning process on a small monolingual dataset, and 3) the inference process of the proposed framework.
Model # of Vocabulary en ja ko Bloomz 250K 26.0 46.5 72.4 Llama-2 32K 29.5 66.3 94.6 GPT-3.5 / 4 100K 25.9 58.6 61.3 Mono-Ja 30K - 28.1 - Mono-Ko 32K - - 32.3
Table 2: Average number of tokens after tokenization. We evaluate tokenization efficiency on the FLoRes-200 dataset for English, Japanese, and Korean Languages. Mono-Ja, and Mono-Ko indicate the tokenizer of pre-trained monolingual model for Japanese and Korean respectively.
3.1 Model Architecture
We illustrate the model architecture of MuMo in Fig. 2.
Pre-trained Multilingual Model
We consider a setting in which a pre-trained multilingual model $f_{\text{multi}}$ is given. The model consists of 1) Transformer layers that consist of attention and feed-forward network, and 2) an output embedding layer called language model (LM) head. We denote $\mathcal{V}_{\text{multi}}$ as the multilingual vocabulary set of the model. objective, as $\mathcal{L}_{\text{MLE}}(p_{\text{multi}},\mathbf{x})=\sum_{t=1}^{T}\log p_{\text{multi}}(x_{t}|\mathbf{x}_{<t})$
Target Monolingual LM Head
The primary concept involves modifying pre-trained representations to predict a singular token unit within a target monolingual vocabulary $\mathcal{V}_{\text{mono}}$. We refer to the Target Monolingual LM head as $f_{\text{mono}}$, which is composed of two main components: a feed-forward network (FFN) and an output linear layer, represented as $g_{\text{mono}}:\mathbb{R}^{d_{{}_{\text{mono}}}}\rightarrow\mathbb{R}^{|\mathcal{V}_{\text{mono}}|}$:
$\text{FFN}(h)=q(W_{1}^{\top}h)W_{2}\in\mathbb{R}^{d_{\text{mono}}}$ (2)
where $W_{1}\in\mathbb{R}^{d_{\text{multi}}\times d_{\text{ffn}}}$ and $W_{2}\in\mathbb{R}^{d_{\text{ffn}}\times d_{\text{mono}}}$ are the weight matrices, $q$ is non-linearity function, and $d_{\text{mono}}$ represents the dimension of the target language representaiton. We set $d_{\text{ffn}}$ as $d_{\text{multi}}/4$, and the non-linearity function $q$ as SwiGLU (Shazeer, 2020). The output linear layer $g_{\text{mono}}$ then generates a word.
$\begin{split}f_{\text{mono}}(h)=g_{\text{mono}}(\text{FFN}(h))\in\mathbb{R}^{|\mathcal{V}_{\text{mono}}|}\end{split}$ (3)
MuMo LM Head
Note that the output space of $f_{\text{mono}}$ is restricted to tokens in the $\mathcal{V}_{\text{mono}}$. Inspired by Lan et al. (2023), we simply extend the $f_{\text{mono}}$ by concatenating the output linear layer of pre-trained multilingual model. This is particularly useful when there is no suitable token in $\mathcal{V}_{\text{mono}}$ to predict, such as special symbols or alphabet-based tokens for non-alphabet languages.
Formally, given context representation $h_{t-1}$, the output of the MuMo LM head is computed as:
$f_{\text{mumo}}(h_{t-1})=$
$[f_{\text{multi}}(h_{t-1});f_{\text{mono}}(h_{t-1})]\in\mathbb{R}^{|\mathcal{V}_{\text{multi}}|+|\mathcal{V}_{\text{mono}}|}$ (4)
where the symbol ; indicates the concatenation of two vectors, and the $f_{\text{mumo}}$ indicates the output of the MuMo LM head. Thus, the MuMo LM head is composed of a combination of the pre-trained language model head and Target Monolingual LM head.
3.2 Fine-tuning
In the proposed framework, we only fine-tune the target monolingual LM head $f_{\text{mono}}$ leveraging a small given target monolingual dataset. Note that the parameters of the pre-trained multilingual model remain frozen during the process. The model is fine-tuned by maximizing the log-likelihood of a sequence:
$\underset{{f_{\text{mono}}}}{\max}~{}\mathcal{L}_{\text{MLE}}(p_{\text{mumo}},\mathbf{x})=\sum_{t=1}^{T}\log p_{\text{mumo}}(x_{t}|\mathbf{x}_{<t})$
where $p_{\text{mumo}}(x_{t}|\mathbf{x}_{<t})=\operatorname{Softmax}(f_{\text{mumo}}(h_{t-1}))$.
3.3 Inference
Despite the availability of direct generation based on the $p_{\text{mumo}}$, the newly initialized Target Monolingual LM head, which is trained on limited data, may be constrained by generalization capabilities beyond the training dataset. The key concept is to leverage the probabilistic knowledge acquired by the pre-trained model $p_{\text{multi}}$, which has been extensively trained on large text corpora.
3.3.1 Step 1: Top-$k$ Selection
Initially, we select top $k$ candidates based on the probability $p_{\text{mumo}}(x_{t}|\mathbf{x}_{<t})$. Given the fact that we do not modify the input embedding of the pre-trained model, we are unable to feed the predicted word if a word does not belong in $\mathcal{V}_{\text{multi}}$ during the subsequent iteration. Instead, we input the predicted word as the tokenization units of the pre-trained vocabulary. For example, let’s consider the Korean word “수소”, which corresponds to a sequence of two tokens (“수”, “소”) in $\mathcal{V}_{\text{multi}}$. If the Korean word “수소” is selected among the Top-$k$ candidates, we employ these two multilingual tokens.
3.3.2 Step 2: Verification
Then, the feasibility of these potential completions is measured using the log-joint probability distribution over $p_{\text{multi}}$. To account for shorter sequences naturally having higher scores (Jean et al., 2015; Murray and Chiang, 2018), we normalize each candidate’s score by its token length.
We measure the feasiblity for a candidate sequence as follows:
$\begin{split}\sigma(\mathbf{c}^{i})&=\frac{1}{l^{i}}\sum_{k=1}^{l^{i}}{\log p_{\text{multi}}(c_{t+k}^{i}|c_{<t+k}^{i},\mathbf{x}_{<t})}\end{split}$ (6)
where $\mathbf{c}^{i}$ symbolizes a predicted token within the top-$k$ candidates, $p_{\text{multi}}$ represents the probability as determined by the pre-trained multilingual model, and $l^{i}$ corresponds to the sequence length of the candidate $\mathbf{c}^{i}$.
From the $k$ candidates, the ultimate prediction can be derived from both deterministic and stochastic manners, depending on decoding strategies.
4 Experiments
4.1 Setup
Languages
We focus on two non-roman alphabetical languages: Korean and Japanese. These languages constitute over 0.05% of the pre-training corpus used in Llama-2 (Touvron et al., 2023). They are of particular interest due to their tokenization disparity when compared to English scripts (Ahia et al., 2023; Petrov et al., 2023). Notably, previous work (Bandarkar et al., 2023) has shown impressive results on natural language understanding tasks for these languages using Llama-2.
Language Language Family Pre-trained Tokenizer Korean Koreanic EleutherAI/polyglot-ko-12.8b Japanese Japonic rinna/japanese-gpt-neox
Table 3: Selected languages and tokenizers. We utilize the tokenizers to construct $\mathcal{V}_{\text{mono}}$ in each language.
Model
We utilize the Llama-2 13B model (Touvron et al., 2023) for all experiments. We observed some language alignment discrepancies between instructions and responses when using the Llama-2 13B chat model.111meta-llama/Llama-2-13b-chat To address the issue, we conduct multilingual instruction tuning (Muennighoff et al., 2022) for English, Korean, and Japanese languages using the ShareGPT and Alpaca (Chen et al., 2023c). This process improve the model’s fluency in each language (Muennighoff et al., 2022; Chen et al., 2023b).
Implementation of MuMo
To construct targeted monolingual vocabularies in MuMo Framework, we levergage the tokenizers from the off-the-shelf model, as shown in Table 3. We selected monolingual tokens by filtering vocabulary items based on the Unicode range of each monolingual script. Additionally, we excluded items from the selection if they were already present in the pre-trained vocabulary.
Regarding the initialization of $g_{\text{mono}}$, we utilize the LM head of the pre-trained multilingual model. For example, when the Korean word "태양" is tokenized into subword units (“\0xed”, …, “\0x91”) using the pre-trained vocabulary, we initialize the Korean LM head of "태양" by taking the mean of the corresponding subword embeddings of the multilingual LM head. This process ensures that the initialized embeddings of Target Monolingual head represent the original word in the multilingual context.
Fine-tuning
We only train the Target Monolingual LM head $g_{\text{mono}}$ with the translated ShareGPT and Alpaca datasets (Chen et al., 2023c) in Korean, and Japanese. The training is done with 1500 steps, using 8 A100-40GB GPUs, with one batch consisting of 128 examples. We use the AdamW (Loshchilov and Hutter, 2019) optimizer with a learning rate of 0.001, weight decay of 0.01, and 150 steps of warm-up.
Summarization (0-shot) Translation (3-shot) Lang Method ROUGE-2 ROUGE-L Tokens/sec Speed Up BLEU Tokens/sec Speed Up Ko Vanilla Decoding 20.7 36.1 28.9 1.00x 21.2 29.8 1.00x Spec. (w/o Rejection) 18.7 33.5 35.2 1.21x 18.6 36.5 1.22x Spec. 20.3 35.2 27.5 0.95x 21.5 29.2 0.98x Shortlisting 20.5 36.3 30.6 1.06x 19.5 32.7 1.03x MuMo (Ours) 20.3 35.9 55.3 1.92x 21.7 50.9 1.70x Ja Vanilla Decoding 11.3 26.6 29.3 1.00x 26.3 33.4 1.00x Spec. (w/o Rejection) 10.8 24.2 35.4 1.21x 22.7 39.9 1.21x Spec. 11.6 26.5 28.5 0.97x 26.0 29.7 1.03x Shortlisting 11.4 26.3 30.3 1.03x 25.2 34.9 1.04x MuMo (Ours) 11.6 26.3 59.2 2.02x 24.3 58.3 1.75x
Table 4: Comparative study of Language Model (LM) inference speed. The column labeled “Speed Up” represents the relative performance improvement in inference speed compared to the vanilla decoding method. The highest performance in each category is highlighted in Boldface, and the second highest score is underlined. All models use sampling-based decoding. MuMo outperforms the compared baselines in the inference speed. Detailed information about the generation hyperparameters, including those used for sampling-based decoding, can be found in Appendix A.
Evaluation
We choose two representative generation tasks: summarization and translation. For summarization, we use 500 examples from XLSum (Hasan et al., 2021), and for translation, we use 500 examples from the FLoRes-200 (Goyal et al., 2022) dataset. We translate English sentences to each target language sentence.
For each task, we report 0-shot results for summarization, and 3-shot results for translation. We set the maximum sequence length as 512. We utilize flash-attention 2 (Dao, 2023) and bfloat16 types for text generation.
Metrics
In the summarization task, we gauge the reliability of the generated content by calculating the ROUGE-2 and ROUGE-L (Lin, 2004) scores, averaging the results across 5 different generated summaries. Likewise, for the translation task, we measure the quality of the translations by computing the BLEU (Papineni et al., 2002) score, again averaging over 5 translation results. [^2] We utilize SacreBLEU scores with the signature BLEU —nrefs:1 —case:mixed —eff:no —tok:ko,ja-mecab—smooth:exp —version:2.2.0. We report Tokens/sec to measure the inference speed of the models.
4.2 Baselines
We consider the following several baselines for the comparison with the proposed method.
Vanilla Decoding
The autoregressive generation is to sequentially sample the subsequent word based on the probability distribution over the pre-trained vocabulary. This approach serves as the standard against which improvements are measured. Accounting for the nature of task, all the baselines and our framework utilizes sampling-based decoding strategy with temperature as 0.1, $k$ as 10 for top-$k$ sampling (Fan et al., 2018) and $p$ as 0.7 for nucleus sampling (Holtzman et al., 2020).
Speculative Decoding
Speculative decoding approach (Chen et al., 2023a; Leviathan et al., 2023) employs a preliminary "draft" model to rapidly generate a set of token candidates at each decoding step. Subsequently, these candidates undergo a validation process by the original language model to ascertain their likelihood as plausible continuations of the text. We implement two variants of this method: one with the capability to reject unsuitable candidates (Spec.) and another without its rejection module (Spec. w/o Rejection). For the draft model, we utilize Llama-2 7B (Touvron et al., 2023). Following the implementation of Chen et al. (2023a), we generate 5 draft tokens at each iteration.
Lexical Shortlisting
Lexical Shortlisting (Shortlisting) (Abdaoui et al., 2020; Ushio et al., 2023), or vocabulary selection, is the approach that optimizes the decoding process by allowing it to generate a word within a set of tokens during the inference stage (Ushio et al., 2023). We implement to filter out tokens that are not present within the corresponding target language subset of the mC4 corpus (Xue et al., 2021), as Ushio et al. (2023).
Summarization (0-shot) Translation (3-shot) Method Update Param. Dataset size (Tokens) ROUGE-2 ROUGE-L Morphemes/sec Speed Up BLEU Morphemes/sec Speed Up Vanilla Fine-tuning 13.0B 365M 20.7 36.1 9.8 1.00x 21.2 10.1 1.00x Vocabulary Expansion 13.1B 365M 13.7 23.1 17.1 1.92x 12.3 20.2 2.00x Vocabulary ${}^{\dagger}$ 13.1B 60B + 365M 20.3 37.3 20.5 2.12x 20.3 23.1 2.29x MuMo (Ours) 70M 365M 20.5 36.3 15.3 1.73x 21.7 17.2 1.71x
Table 5: Comparsion with the fine-tuning strategies. The column labeled “Speed Up” represents the relative performance improvement in inference speed compared to Vanilla Fine-tuning. Vocabulary ${}^{\dagger}$ was pre-trained on over 60B tokens, comprised of both Korean and English text corpora. Other methods are only trained with the instruction dataset (365M) (Chen et al., 2023c), ShareGPT and Alpaca translated in Korean. The Boldface signifies the superior performances, and the second highest score is underlined.
4.3 Results
Table 4 shows the generation results in both summarization and translation tasks. For the summarization task in Korean, MuMo outperforms all baselines in terms of speed, achieving a 1.92x speed-up over the Vanilla Decoding while maintaining competitive ROUGE scores. In translation, MuMo again demonstrates superior efficiency with a 1.70x speed-up and even shows an improvement in BLEU score compared to Vanilla Decoding.
In the case of Japanese, the results are similar, with MuMo achieving a 2.02x speed-up in summarization and a 1.75x speed-up in translation. The ROUGE and BLEU scores for MuMo are on par with or slightly below Vanilla Decoding, indicating that the increase in speed does not significantly compromise the quality of the output.
Shortlisting shows only marginal gains in speed across both languages and every tasks, while preserving the generation capability. This is likely because the relative computational cost of processing the embedding matrix is reduced in larger models, making vocabulary reduction less impactful (Berard et al., 2021; Ushio et al., 2023). On the other hand, the Spec. heavily relies on the capacity of the draft model, as shown as the comparison with (Spec. w/o Rejection). If the draft model lacks of sufficient multilingual capacity, it may not generate high-quality candidates, leading to a lower acceptance rate by the original model and thus reduced efficiency.
The superior performance of MuMo in terms of inference speed can be primarily attributed to its capability to predict larger linguistic units compared to those in the pre-trained vocabulary. We found that the target language tokens in $\mathcal{V}_{\text{mono}}$ are typically tokenized into 3-4 separate tokens in $\mathcal{V}_{\text{multi}}$, suggesting that the decoding step could potentially be reduced by 3-4 times. It is hypothesized that the inference speed is significantly influenced by the disparity between the pre-trained multilingual vocabulary and the target language.
5 Further Analysis
Summarization (0-shot) Translation (3-shot) LM Head Initialization ROUGE-2 ROUGE-L BLEU Mono-init 20.7 36.2 21.5 Random-init 19.2 35.5 17.2 Multi-Init 20.3 36.3 21.7
Table 6: Comparative analysis for the initialization strategy. Mono-Init signifies to leverage the pre-existing embedding representation. We use the language model head of the monolingual model from EleutherAI/polyglot-ko-12.8b. In the case of Random-Init, we randomly initialize with Gaussian distribution. Multi-Init indicates to leverage multilingual model representation by averaging its subword embedding as the main experiment. The Boldface signifies the superior performances.
Summarization (0-shot) Translation (3-shot) Lang Method ROUGE-2 ROUGE-L Tokens/sec BLEU Tokens/sec Ko MuMo 20.3 35.9 55.3 21.7 50.9 w/o Verification 11.0(-9/3) 26.4(-9.5) 60.8(+5.5) 16.3(-5.4) 62.3(+11.4) Ja MuMo 11.6 26.3 59.2 24.3 58.3 w/o Verification 6.7(-4.9) 20.4(-5.9) 69.1(+9.9) 10.8(-13.5) 73.6(+15.3)
Table 7: Ablation Study. While the exclusion of the verification accelerates approximately 1.2 times in inference speed, it significantly compromises the quality of the generation.
5.1 Comparative Analysis of Fine-Tuning Strategies
In the section, we provide a comparative analysis of three distinct fine-tuning strategies for multilingual models. This analysis aims to highlight the advantages and disadvantages of each strategy, particularly in terms of dataset requirements and the number of parameters to train.
5.1.1 Setup
The two strategies compared in the analysis are:
1. Vanilla Fine-tuning: This strategy, which serves as a baseline, involves fine-tuning a standard multilingual model on the instruction dataset without any modifications to the pre-trained vocabulary.
2. Vocabulary Expansion: Inspired by prior work Chau et al. (2020); Cui et al. (2023), this strategy involves expanding the vocabulary of the pre-trained multilingual model and fine-tuning on the instruction dataset. This method, unlike MuMo, expands not only the LM head but also the token embedding in the input layer. Two implementations of this strategy are considered. The first involves pre-training on large-scale text corpora (60B tokens) [^3] We use the off-the-shelf checkpoint from beomi/llama-2-koen-13b before fine-tuning on the instruction dataset. This strategy is marked with a dagger in Table 5. The second only undergoes the fine-tuning phase on the instruction dataset.
To account for the variability of token unit between the different strategies, we report the inference speed with the morphemes per second (Morphemes/sec), providing a standardized measurement.444python-mecab-ko.
5.1.2 Discussion
Table 5 reveals a consistent trend across both summarization and translation tasks. The vocabulary expansion strategies, which expand the dimension of both the token embeddings and LM head, exhibit significant increases in inference speed, but this is accompanied by a substantial decrease in the quality of the generated output when not trained on large-scale text corpora. This indicates that merely fine-tuning with an expanded vocabulary on a limited downstream dataset may not suffice to maintain high-quality text generation, as suggested by Conneau et al. (2020). Furthermore, while vocabulary expansion with pre-training achieves notable speed improvements, it does not exhibit significant enhancements in generation quality.
In contrast, our proposed method exhibits a modest increase in speed while also slightly improving BLEU scores relative to vanilla fine-tuning. The principal advantage of our method lies in its capacity to attain these results without necessitating vast monolingual text corpora. This approach not only reduces the number of parameters that need to be fine-tuned, making it more parameter-efficient but also lessens the dependency on large-scale data for pre-training, making it a more data-efficient solution.
5.2 Initialization of Target Monolingual LM Head
We investigate the impact of 3 different initialization strategies on the target monolingual LM head $g_{\text{mono}}$ in the Target Monolingual LM head. The first strategy involves leveraging embeddings that correspond to the pre-trained representation of a targeted monolingual LM head, termed as Mono-init. The second strategy is initializing the parameters with random value using Gaussian distribution (Random-init). Lastly, we utilize the embeddings from the pre-trained multilingual LM head (Multi-init), as the main experiment. This is achieved by averaging the output embeddings of the multilingual model.
Table 6 shows that Multi-init achieves a ROUGE-L score of 36.3 and a BLEU score of 21.7, which are close to the 36.2 ROUGE-L and 20.9 BLEU scores of Mono-init. On the other hand, Random-init shows a decrease in performance, with a ROUGE-L score of 35.5 and a BLEU score of 17.2.
The result demonstrates that the Multi-init approach is almost equally effective with Mono-init. This suggests that our framework can be utilized even without a pre-existing target monolingual model, making it applicable to low-resource languages.
5.3 Effectiveness of Verification Step
We design an ablation study to investigate the role of the verification step in the inference process ( Sec. 3.3.2). To assess the impact of the verification step, we generated sequences without employing the verification step.
From the results in Table 7, conducted in both Korean and Japanese, we notice that the overall generation speed is approximately 1.2 times faster when the verification is excluded. However, it is crucial to highlight that the exclusion of the verification step in the inference phase leads to a significant reduction in the generation quality. This is evident in the decrease in ROUGE-2, ROUGE-L, and BLEU scores for both languages when the verification module is not used, as shown in the table. This suggests that while the verification step may slightly slow down the generation process, it plays a vital role in preserving the model’s generation capability.
6 Related Work
Tokenization Disparity
Subword tokenization, a common approach in LMs, is typically data-driven. Most of pre-trained tokenizers, which are often trained on predominantly English corpora, frequently result in excessive fragmentation of non-English scripts (Rust et al., 2021; Zhang et al., 2022a). Ahia et al. (2023); Petrov et al. (2023) have found significant tokenization disparities across languages across popular LLMs (Xue et al., 2021, 2022; Scao et al., 2022; OpenAI, 2023) with a depth analysis. Our work endeavors to address the slowdown in inference that arises due to tokenization disparity in non-alphabetic languages.
Modifying Pre-trained Vocabulary
Previous works have explored the adaptation of pre-trained vocabularies or the addition of new tokens (Artetxe et al., 2020; Rust et al., 2021; Liu et al., 2023), these methods often necessitate extensive pre-training to integrate the new tokens effectively (Wang et al., 2019b; Chau et al., 2020; Cui et al., 2023; Liu et al., 2023). In contrast, our MuMo framework sidesteps the need for fine-tuning the parameters of pre-trained models to preserve the original capabilities of the pre-trained language model. Efforts to select items of pre-trained embedding matrix have been made (Abdaoui et al., 2020; Domhan et al., 2022; Ushio et al., 2023), but these have not yielded significant speed up in our experiment, where the relative size of the embedding layer is small (Bogoychev et al., 2023).
Accelerating LLM Inference
The quest to accelerate inference in auto-regressive large language models (LLMs) has led to a variety of approaches. There has been a proliferation of systems specifically engineered for LLM inference, such as Orca (Yu et al., 2022), FlexGen (Sheng et al., 2023), and StreamingLLM (Xiao et al., 2023). Our proposed methodology can be harmonically integrated with the aforementioned techniques. Speculative decoding (Leviathan et al., 2023; Chen et al., 2023a) have also been explored to increase inference velocity. However, the approach often relies on the assumption that a small model can maintain high fidelity when generating a series of multiple tokens. Moreever, acquiring a small yet competitive model may be tricky, especially in a multilingual setup (Conneau et al., 2020; Bandarkar et al., 2023). Our work distinguishes itself by specifically solving the inference inefficiency that arises from excessive fragmentation in the non-alphabetic context.
Parameter Efficient Cross-lingual Transfer Learning
The curse of multilinguality, which refers a trade-off between the language coverage and model capacity (Conneau et al., 2020), is a significant issue even in massively multilingual models, such as mBERT, XLM-R, and mT5 (Devlin et al., 2019; Conneau et al., 2020; Xue et al., 2021; Ansell et al., 2021). The problem has been mitigated through modular parameter-efficient adaptations of the multilingual models through lightweight adapters (Houlsby et al., 2019): additional trainable parameters inserted into the transformer layers of model (Pfeiffer et al., 2020; Üstün et al., 2020; Vidoni et al., 2020; Parović et al., 2022) for a target language. Previous studies focus on improving the representational capability of multilingual models, rather than directly addressing the issue of speeding up multilingual models. In contrast, our research primarily aims to accelerate the inference for multilingual models in specific target languages.
7 Conclusion
Our study has successfully tackled the challenges in generating text for non-alphabet languages, particularly those associated with excessive fragmentation issues. The approach not only speeds up text generation but also paves the way for more efficient multilingual language applications. Our future work will broaden our experimental scope to languages that were not substantially represented in the pre-trained corpus.
Acknowledgements
We thank Chaehun Park for their valuable discussions and feedback on this paper.
Appendix
Appendix A Implementation Details
Hyperparameter Value temperature 0.1 sampling True $p$ for top-$p$ sampling 0.7 repetition penalty 1.05 exponential decay length penalty (256, 1.03) max sequence length 512 $k$ for top-$k$ sampling 20
Table 8: Hyperparameters used for generation.